OpenStudy (anonymous):

If a baseball is dropped from a helicopter 625 feet above ground, then its distance in feet form the ground t seconds later is modeled by -16t^2+625 How long after the ball is dropped will it hit the ground? 1.56 secs 6.25 secs 9 secs 39.06 secs help please

OpenStudy (anonymous):

When it hits the ground, the distance from the ground will be 0, so you have -16t^2 + 625 = 0. Can you solve for t from there?

OpenStudy (anonymous):

so would you write it as -16^2=-625

OpenStudy (anonymous):

And you can try to see if 625 is a square of anything. It is the square of 25 right?

OpenStudy (anonymous):

yea.

OpenStudy (anonymous):

@seraphic_topaz

OpenStudy (anonymous):

And 16 is a square of 4 right? So from the equation above you can arrive at (4t)^2 = 25^2 And then 4t = 25; t is time and cannot be negative. Can you solve for t from there?

OpenStudy (anonymous):

6.25?

OpenStudy (anonymous):

i thought you had to plug in the 60 secs. like 4(60)=25

OpenStudy (anonymous):

Why do you have to plug in the 60 secs?

OpenStudy (anonymous):

since t=time i always thought you did that

OpenStudy (anonymous):

Oh, you only need to look at the question for specification, like in this question it says "then its distance in feet form the ground *t seconds* later is modeled by -16t^2+625." Everything is in seconds so we don't need to worry about the conversion.

OpenStudy (anonymous):

oh alright i think i got it thank you :)